Link

Documentation automatically generated by TEX-Reporter written by Bert Schönwälder, 1996 July 14, 1996

1 Class: Link

Super class: Object Category: Collections-Support Instance variables: nextLink Comment: Class Link represents a simple record of a pointer to another

Link.

Instance Variables:

nextLink <Link> a pointer referencing the next Link in the chain

1.1 Protocol: accessing

1.1.1 Method nextLink nextLink

"Answer the Link to which the receiver points."

^nextLink 1.1.2 Method nextLink: nextLink: aLink

"Store the argument as the Link to which the receiver refers." ^nextLink := aLink

2 Metaclass: Link class

2.1 Protocol: instance creation

2.1.1 Method nextLink:

nextLink: aLink

"Answer a new instance of the receiver referring to the argument, aLink."

^self new nextLink: aLink; yourself